OpenStudy (anonymous):

My new TV has a perimeter of 164 inches. Its width is 8 inches less than twice the height. How many inches wide is the TV?

OpenStudy (therealmeeeee):

let me tell u how to do this

OpenStudy (anonymous):

ok

OpenStudy (therealmeeeee):

2w+2h = p 2w+2h=164 so w is(=) 2h-8 plug in for w 2(2h-8) + 2h = 164 solve for h 4h-16+2h=164 6h-16 = 164 6h=180 hence h=30 2 times h - 8 is 52 check 52 *2 + 30 *2= 164

OpenStudy (therealmeeeee):

there ya go

OpenStudy (anonymous):

thank you

OpenStudy (therealmeeeee):

give me a medal

OpenStudy (anonymous):

How do I do that?

OpenStudy (therealmeeeee):

click best response

OpenStudy (anonymous):

just did that

OpenStudy (therealmeeeee):

ty

OpenStudy (anonymous):

yw. Just posted another question. Can you help me??